OpenStudy (anonymous):

prove that cos(2s)=2cos^2s - 1

OpenStudy (anonymous):

\[\cos (s+s) = \cos(s).\cos(s) - \sin(s) . \sin(s) = \cos^2(s) - \sin^2 (s)\]

OpenStudy (anonymous):

\[\sin^2 (s) + \cos^2 (s) = 1 \]

OpenStudy (anonymous):

\[\sin^2 (s) = 1- \cos^2 (s)\]

OpenStudy (anonymous):

\[\cos^2 (s) - ( 1 - \cos^2(s) ) = 2\cos^2 (s) - 1\]

OpenStudy (anonymous):

how about the identity 1-2sin^2s?

OpenStudy (anonymous):

\[\cos^2 (s) = 1-\sin^2(s)\]

OpenStudy (anonymous):

\[1 - \sin^2 (s) - \sin^2(s) = 1- 2\sin^2 (s)\]

OpenStudy (anonymous):

ahhh i see Im beginning to get it